OpenStudy (anonymous):

Solve the triangle (c, A,B) a=5 b=6 C=70degrees

OpenStudy (chaise):

Are you familiar with the Cosine rule? \[a^2=b^2+c^2-2bc CosA\]

OpenStudy (callisto):

You can use sine law and cosine law to solve the triangle. As you have 2 sides with an included angle, you can apply cosine law first. c^2 = a^2 +b^2 -2abcosC = 5^2 + 6^2 - 2(5)(6)cos 70 c = 6.36 Then, you can apply sine law to get the rest of the unknowns sinA sinC a sinC 5sin70 ---- = ----- => sinA = ------ = ------ => A = 47.6 a c c 6.36 Can you find out the last angle (angle B) ?

OpenStudy (callisto):

You can use the above method to find angle B, or use angle sum of triangle to find it. Note that angle sum of triangle is angle A + angle B + angle C =180 degrees